OpenStudy (anonymous):

determine the measure of 0 in radian and degree measures. so length = 8cm, and radius= 5 cm

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

Given the radius 5 cm, what's the circumference?

OpenStudy (anonymous):

um im not sure, would it be 15.75?

jimthompson5910 (jim_thompson5910):

no, it's pi times diameter (not pi times radius)

OpenStudy (anonymous):

so 31.5 ?

jimthompson5910 (jim_thompson5910):

yes or just 10pi

jimthompson5910 (jim_thompson5910):

what fraction of the circle is 8 cm (compared to 10pi or 31.5 cm)?

OpenStudy (anonymous):

um quater? so 31.5/8? 3.93?

jimthompson5910 (jim_thompson5910):

you have it flipped the wrong way

jimthompson5910 (jim_thompson5910):

it should be 8/31.5 = 0.25396825, which is roughly a quarter, yes

OpenStudy (anonymous):

okay and then what shall I do?

jimthompson5910 (jim_thompson5910):

so what you do is you multiply 0.25396825 with 2pi (since one full revolution is 2pi radians) so the angle in radians is 0.25396825 * 2pi = 1.595729577 radians (this is an approximation)

jimthompson5910 (jim_thompson5910):

It turns out that you can also use the idea that L = theta*r where L is the arc length and theta is the angle in radians, so... L = theta*r 8 = theta*5 8/5 = theta theta = 8/5 theta = 1.6 radians (which is the exact angle in radians)

jimthompson5910 (jim_thompson5910):

notice how 1.595729577 is very very close to 1.6

OpenStudy (anonymous):

oh yes yes and then how would I get the degree meeasure?

jimthompson5910 (jim_thompson5910):

multiply this with 180/pi to convert to degrees

jimthompson5910 (jim_thompson5910):

Degree Measure = (1.6)*(180/pi) = ???

OpenStudy (anonymous):

91.7 :) wow thank you so much !

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):